# Partially Retargeted Balancing Weights for Causal Effect Estimation Under Positivity Violations

### v1

@@ Discussion
Positivity violations create substantial challenges when estimating causal effects with observational data. These challenges necessitate researchers to confront trade-offs between estimator bias, variance, and maintaining the original target population. We have proposed a novel weighting procedure that achieves reduced estimator bias and variance through modifying the target population for only a subset of covariates. In doing so, our proposed weights can be derived under an relaxed positivity assumption. We have shown that our proposed estimator 1) is consistent for the original estimand when either the implied propensity score model is correct or the set of treatment effect modifiers is properly specified; and 2) is consistent for a slightly modified estimand that is simple to characterize when these conditions do not hold. Furthermore, we have shown that our proposed estimator achieves reduced asymptotic variance under the relaxed positivity assumption. We have proposed an algorithm for identifying the set of covariates not balanced to the target population; the corresponding design- and model-based estimators perform well across applications to synthetic data, EHR data, and when transporting RCT effects. Thus, our methods allow analysts to derive weights that result in estimators with minimal bias and variance with respect to a meaningful and interpretable target population.
@@
While we focus on estimating the ATE, our proposed weighting procedure can be readily extended to other causal effect estimation problems as discussed in Section 5 and supplementary Section 3. Future work includes exploring the theoretical properties of these extensions. Our proposed extension to distributional balancing weights may be especially promising as these weights do not require strong parametric assumptions on the outcome models. In addition, there is the potential for combining the intuition behind our proposed method and modified treatment policies to obtain estimators with reduced bias and variance that correspond to both a treatment and population of interest. While our weighting procedure does perform well when there are positivity violations with higher dimensional covariates (e.g, $p=100$), there is a need for additional methods that address the combined challenges of variable selection (in the context of unmeasured confounding) and positivity violations in $p>>n$ scenarios. Further, while the infeasibility of Problem (3) is a clear indication that our proposed method or minimal weights need to be used to derive weights, it is generally unclear when estimator performance may be impacted by the overlap levels for a given research problem. Thus, future work includes developing measures of overlap that are directly related the impact of overlap on estimator MSE.
@@
$Q_{2K+L\times 2n}$ $=\begin{bmatrix}c_{1}(\bm{X}_{1})&c_{1}(\bm{X}_{2})&\cdots&c_{1}(\bm{X}_{n})&0&0&\cdots&0\\ c_{K}(\bm{X}_{1})&c_{K}(\bm{X}_{2})&\cdots&c_{K}(\bm{X}_{n})&0&0&\cdots&0\\ 0&0&\cdots&0&c_{1}(\bm{X}_{1})&c_{1}(\bm{X}_{2})&\cdots&c_{1}(\bm{X}_{n})&\\ 0&0&\cdots&0&c_{K}(\bm{X}_{1})&c_{K}(\bm{X}_{2})&\cdots&c_{K}(\bm{X}_{n})&\\ g_{1}(\bm{X}_{1})&g_{1}(\bm{X}_{2})&\cdots&g_{1}(\bm{X}_{n})&-g_{1}(\bm{X}_{1})&-g_{1}(\bm{X}_{2})&\cdots&-g_{1}(\bm{X}_{n})\\ g_{L}(\bm{X}_{1})&g_{L}(\bm{X}_{2})&\cdots&g_{L}(\bm{X}_{n})&-g_{L}(\bm{X}_{1})&-g_{L}(\bm{X}_{2})&\cdots&-g_{L}(\bm{X}_{n})\\ \end{bmatrix}$
This problem is now in the form of Tseng and Bertsekas, (1991). The dual of this problem is,
@@
subject to $\>\>\>\text{no constraint on}\>\>\lambda$
where $q(\lambda)=\sum_{j=1}^{2n}h^{*}_{j}(Q_{j}^{T}\lambda)$ and $h^{*}_{j}(t)=\sup_{s_{j}}\{ts_{j}-h(s_{j})\}$. Consider $j<n+1$. Then we have that,
@@
$=t-tZ_{j}w_{j}^{*}-Z_{j}h(1-w_{j}^{*})-(1-Z_{j})h(1)$
Next we have that,
@@
  $\implies\>\>1-(h^{\prime})^{-1}(t)$ $=w^{*}_{j}$
Then we have that,
@@
$=-(1-Z_{j})\{t-t(h^{\prime})^{-1}(t)+h[(h^{\prime})^{-1}(t)]-h(1)\}-h(1)+t$
Let $\rho(t)=t-t(h^{\prime})^{-1}(t)+h[(h^{\prime})^{-1}(t)]-h(1)$. This gives,
@@
subject to $\>\>\>\text{no constraint on}\>\>\lambda$
@@
$=\frac{1}{n}\sum_{j=1}^{n}-Z_{j}\rho(Q_{j}^{T}\lambda)-(1-Z_{j})\rho(Q_{j+n}^{T}\lambda)+Q_{j}^{T}\lambda+Q_{j+n}^{T}\lambda$
@@
$w^{*}_{j}$ $=\rho^{\prime}\{B^{-}(\bm{X}_{j})^{T}\hat{\lambda}\}\>\>\text{for}\>\>Z_{j}=0\>\>;\>\>j=1,\ldots n$
where $\hat{\lambda}$ is the solution to the dual problem. Let $\lambda=\begin{bmatrix}\bm{\alpha}_{0_{K\times 1}}\\ \bm{\alpha}_{1_{K\times 1}}\\ \bm{\gamma}_{L\times 1}\end{bmatrix}$ and let $c(\bm{X}_{j})=\{c_{1}(\bm{X}_{j}),\ldots c_{K}(\bm{X}_{j})\}$ and $g(\bm{X}_{j})=\{g_{1}(\bm{X}_{j}),\ldots,g_{L}(\bm{X}_{j})\}$. Then the dual formulation simplifies to:
@@
subject to $\>\>\>\text{no constraint on}\>\>\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma}$
@@
$+\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X}_{i})+\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X}_{i})$
@@
$\times P(\exists\bm{X}_{j}\in R_{j}(\bm{x}^{*}),Z_{j}=1,j=1,\ldots,3^{p})$
@@
Suppose Assumptions Assumption 1: Strong ignorability, either Assumption 2: Positivity or Assumption 2: Relaxed positivity, and Assumption 3: Balance function regularity conditions hold and $\rho$ has a continuous first derivative. Let $\bm{\theta}=(\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma}).$ Then,
@@
We use similar proof techniques to the proof of proposition 1 in Källberg and Waernbaum, (2023). First, we prove the following lemma.
@@
$H_{k_{1},k_{2}}(\bm{\theta})$ $=E[-e(\bm{X})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X})+\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})c_{k_{2}}(\bm{X})$
$-\{1-e(\bm{X})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X})-\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})c_{k_{2}}(\bm{X})],$
$H_{l_{1},l_{2}}(\bm{\theta})$ $=E[-e(\bm{X})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X})+\bm{\gamma}^{T}g(\bm{X})\}g_{l_{1}}(\bm{X})g_{l_{2}}(\bm{X})$
$-\{1-e(\bm{X})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X})\}g_{l_{1}}(\bm{X})g_{l_{2}}(\bm{X})],$
$H_{k_{1},l_{1}}(\bm{\theta})$ $=E[-e(\bm{X})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})g_{l_{1}}(\bm{X})$
$-\{1-e(\bm{X})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})g_{l_{1}}(\bm{X})].$
We can construct an estimator for $H(\bm{\theta})$, $\hat{H}(\bm{\theta})$ using the sample equivalents of $H_{k_{1},k_{2}}(\bm{\theta}),H_{l_{1},l_{2}}(\bm{\theta}),$ and $H_{k_{1},l_{1}}(\bm{\theta})$. Let $c_{i}=-e(\bm{X}_{i})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X}_{i})\}+-\{1-e(\bm{X_{i}})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X_{i}})\}$. By the strict concavity of $\rho$, $c_{i}>0$. Let $v^{*}=[\sqrt{c_{1}},\ldots,\sqrt{c_{n}}]$. Let $v$ be any vector of length $n$. Then $v^{T}\hat{H}(\bm{\theta})v=v^{T}\odot v^{*^{T}}B(\bm{X})^{T}B(\bm{X})v\odot v^{*}>0$ since $B(\bm{X})$ is full rank by Assumption Assumption 3: Balance function regularity conditions. Then $\hat{H}$ is positive definite and since $E[\hat{H}(\bm{\theta})]=H(\bm{\theta})$, $\hat{H}(\bm{\theta})\to_{p}H(\bm{\theta})$, $H(\bm{\theta})$ is positive definite and $l(\bm{\theta})$ is strictly convex as desired. ∎
@@
$\tau$ $=E[\frac{Z_{i}Y_{i}}{e(\bm{X}_{i})}]-E[\frac{(1-Z_{i})Y_{i}}{1-e(\bm{X}_{i})}]$
@@
Then by Proposition 5 and the law of large numbers for averages with estimated parameters (Boos and Stefanski, (2013) Theorem 7.3),
@@
Assume $\mu_{1}(\bm{x})=\bm{\beta}_{1}^{T}c(\bm{x})+\bm{\lambda}^{T}g(\bm{x})$ and $\mu_{0}(\bm{x})=\bm{\beta}_{0}^{T}c(\bm{x})+\bm{\lambda}^{T}g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}\in\mathbb{R}^{L}$. We will use a similar proof technique by Zhao and Percival, (2017) to show that augmenting the our proposed estimator with estimated outcome regressions does not change the estimator. Let $\hat{\mu}_{0}(\bm{x})$, $\hat{\mu}_{1}(\bm{x})$, and $\hat{e}(\bm{x})$ be outcome regression and propensity score estimates. Then the standard doubly robust estimator for the ATE is,
@@
$+\sum_{Z_{i}=0}w^{*}(\bm{X}_{i})[\hat{\bm{\beta}}_{0}^{T}c(\bm{X}_{i})+\hat{\bm{\lambda}}^{T}g(\bm{X})_{i}]-\frac{1}{n}\sum_{i}\hat{\bm{\beta}}_{0}^{T}c(\bm{X}_{i})+\hat{\bm{\lambda}}^{T}g(\bm{X})_{i}$
@@
$-\hat{\bm{\beta}}_{0}^{T}[\sum_{i}c(\bm{X}_{i})-\sum_{i}w^{*}(\bm{X}_{i})(1-Z_{i})c(\bm{X}_{i})]$
$+\hat{\bm{\lambda}}^{T}[\sum_{i}w^{*}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})-\sum_{i}w^{*}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})]$
@@
and $\hat{\tau}_{DR}=\hat{\tau}_{w^{*}}$. Since $\hat{\tau}_{DR}\to_{p}\tau$ we also have $\hat{\tau}_{w^{*}}\to_{p}\tau$. ∎
@@
Assume $\mu_{1}(\bm{x})=\bm{\beta}_{1}^{T}c(\bm{x})+\bm{\lambda}_{1}^{T}g(\bm{x})$ and $\mu_{0}(\bm{x})=\bm{\beta}_{0}^{T}c(\bm{x})+\bm{\lambda}_{0}^{T}g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}_{z}\in\mathbb{R}^{L}$. Then by Proposition 3 and the law of large numbers for averages with estimated parameters (Boos and Stefanski, (2013) Theorem 7.3)
@@
$-\frac{1}{n}\sum_{i=1}^{n}w^{*}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}-\mu_{0}(\bm{X}_{i})\}+\frac{1}{n}\sum_{i=1}^{n}\{w^{*}_{0}(\bm{X}_{i})(1-Z_{i})-1\}\mu_{0}(\bm{X}_{i})$
$+\frac{1}{n}\sum_{i=1}^{n}\tau(\bm{X}_{i})-\tau$
$=\frac{1}{n}\sum_{i=1}^{n}S_{i}+R_{1}+R_{2}$
where,
@@
$+\sum_{i=1}^{n}\{w_{1}^{*}(\bm{X}_{i})Z_{i}-w_{0}^{*}(\bm{X}_{i})(1-Z_{i})\}\bm{\lambda}^{T}g(\bm{X}_{i})$
Note that $R_{2}=0$ by the constraints in the optimization problem. Then, since $S_{i}$ takes the same form as the efficient score for the ATE in Hahn, (1998), $\hat{\tau}_{w^{*}}$ is asymptotically normal and semiparametrically efficient as long as $R_{1}$ is $o_{p}(n^{-1/2}).$
We will focus on the first component of $R_{1}$,
$|R_{1}^{1}|$ $=|\frac{1}{n}\sum_{i=1}^{n}(w^{*}_{1}(\bm{X}_{i})-\frac{1}{e(\bm{X}_{i})})Z_{i}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}|$
$=|\int(w^{*}_{1}(\bm{X}_{i})-\frac{1}{e(\bm{X}_{i})})\{Z_{i}Y_{i}-Z_{i}\mu_{1}(\bm{X}_{i})\}dF_{n}(\bm{x})|$
[…]
$=o_{p}(1)O_{p}(n^{-1/2})=o_{p}(n^{-1/2})$
@@
$\text{Var}(S_{i}+R_{1i})=\text{Var}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}-\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}+\{\tau(\bm{X}_{i})-\tau\}]$
@@
$+\text{Var}\{\tau(\bm{X}_{i})-\tau\}-2\text{Cov}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\},\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}]$
$+2\text{Cov}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\},\{\tau(\bm{X}_{i})-\tau\}]$
$-2\text{Cov}[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\},\{\tau(\bm{X}_{i})-\tau\}]$
We have that,
@@
$=-2E[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}]\times E[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}]]$
@@
$\times E[\tilde{w}_{0}(\bm{X}_{i})\{1-e(\bm{X}_{i})\}\{\mu_{0}(\bm{X}_{i})-\mu_{0}(\bm{X}_{i})\}]$
$=0\times 0=0$

### v2

@@ Discussion
Positivity violations create substantial challenges when estimating causal effects with observational data. These challenges necessitate researchers to confront the trade-offs between estimator bias, variance, and maintaining the original target population of interest. In this work, we proposed a novel weighting procedure that achieves reduced estimator bias and variance through relaxing the balancing constraints to the target population. In doing so, our proposed weights can be derived under an relaxed positivity assumption. Our approach offers a choice in between direct balancing and overlap weights (which are special cases of our method), allowing the analyst to navigate between preservation of the original estimand and mitigation of positivity violations. We have shown that our proposed estimator 1) is consistent for the original estimand when either the implied propensity score model is correct or the set of treatment effect modifiers is properly specified; and 2) is consistent for a slightly modified estimand that is simple to characterize when these conditions do not hold. Furthermore, we have shown that our proposed estimator achieves reduced asymptotic variance under the relaxed positivity assumption. To aid in the practical use of our method, we proposed an algorithm for identifying the set of covariates balanced to a modified population; the corresponding design- and model-based estimators perform well across applications to synthetic data, EHR data, and when transporting RCT effects.
@@
While we focus on estimating the ATE, our proposed weighting procedure can be readily extended to other causal effect estimation problems as discussed in Section 3.3 and supplementary Section 3. Our proposed extension to distributional balancing weights may be especially promising as these weights do not require strong parametric assumptions on the outcome models. Future work includes exploring the theoretical properties of these extensions. There is also potential for combining the intuition behind our proposed method and modified treatment policies to obtain estimators with reduced bias and variance that correspond to both a treatment and population of interest. While our weighting procedure does perform well when there are positivity violations and higher dimensional covariates (e.g, $p=100$), there is a need for additional methods that address the combined challenges of variable selection (in the context of unmeasured confounding) and positivity violations in $p>>n$ scenarios. Further, while the infeasibility of Problem (3) is a clear indication that our proposed method or minimal weights need to be used to derive weights, it is generally unclear when estimator performance may be impacted by overlap levels for a given research problem. Thus, future work includes developing measures of overlap that are directly related the impact of overlap on estimator MSE.
Funding
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2237827.
Disclosure Statement
The authors report there are no competing interests to declare.
Data Availability Statement
The data that supports Section 7 are 1) MIMIC-III data: available with queries provided at https://github.com/MIT-LCP/mimic-code; 2) RCT data: available at https://doi.org/10.7910/DVN/ZJVVQZ; and 3) Health system target population data: not publicly available.
@@
$Q_{2K+L\times 2n}$ $=\begin{bmatrix}c_{1}(\bm{X}_{1})&c_{1}(\bm{X}_{2})&\cdots&c_{1}(\bm{X}_{n})&0&0&\cdots&0\\ c_{K}(\bm{X}_{1})&c_{K}(\bm{X}_{2})&\cdots&c_{K}(\bm{X}_{n})&0&0&\cdots&0\\ 0&0&\cdots&0&c_{1}(\bm{X}_{1})&c_{1}(\bm{X}_{2})&\cdots&c_{1}(\bm{X}_{n})&\\ 0&0&\cdots&0&c_{K}(\bm{X}_{1})&c_{K}(\bm{X}_{2})&\cdots&c_{K}(\bm{X}_{n})&\\ g_{1}(\bm{X}_{1})&g_{1}(\bm{X}_{2})&\cdots&g_{1}(\bm{X}_{n})&-g_{1}(\bm{X}_{1})&-g_{1}(\bm{X}_{2})&\cdots&-g_{1}(\bm{X}_{n})\\ g_{L}(\bm{X}_{1})&g_{L}(\bm{X}_{2})&\cdots&g_{L}(\bm{X}_{n})&-g_{L}(\bm{X}_{1})&-g_{L}(\bm{X}_{2})&\cdots&-g_{L}(\bm{X}_{n})\\ \end{bmatrix}$.
This problem is now in the form of Tseng and Bertsekas (1991). The dual of this problem is
@@
subject to $\>\>\>\text{no constraint on}\>\>\lambda,$
where $q(\lambda)=\sum_{j=1}^{2n}h^{*}_{j}(Q_{j}^{T}\lambda)$ and $h^{*}_{j}(t)=\sup_{s_{j}}\{ts_{j}-h(s_{j})\}$. Consider $j<n+1$. Then we have that
@@
$=t-tZ_{j}w_{j}^{*}-Z_{j}h(1-w_{j}^{*})-(1-Z_{j})h(1).$
Next we have that
@@
  $\implies\>\>1-(h^{\prime})^{-1}(t)$ $=w^{*}_{j}.$
Then we have that
@@
$=-(1-Z_{j})\{t-t(h^{\prime})^{-1}(t)+h[(h^{\prime})^{-1}(t)]-h(1)\}-h(1)+t.$
Let $\rho(t)=t-t(h^{\prime})^{-1}(t)+h[(h^{\prime})^{-1}(t)]-h(1)$. This gives
@@
subject to $\>\>\>\text{no constraint on}\>\>\lambda,$
@@
$=\frac{1}{n}\sum_{j=1}^{n}-Z_{j}\rho(Q_{j}^{T}\lambda)-(1-Z_{j})\rho(Q_{j+n}^{T}\lambda)+Q_{j}^{T}\lambda+Q_{j+n}^{T}\lambda.$
@@
$w^{*}_{j}$ $=\rho^{\prime}\{B^{-}(\bm{X}_{j})^{T}\hat{\lambda}\}\>\>\text{for}\>\>Z_{j}=0\>\>;\>\>j=1,\ldots n,$
where $\hat{\lambda}$ is the solution to the dual problem. Let $\lambda=\begin{bmatrix}\bm{\alpha}_{0_{K\times 1}}\\ \bm{\alpha}_{1_{K\times 1}}\\ \bm{\gamma}_{L\times 1}\end{bmatrix}$ and let $c(\bm{X}_{j})=\{c_{1}(\bm{X}_{j}),\ldots c_{K}(\bm{X}_{j})\}$ and $g(\bm{X}_{j})=\{g_{1}(\bm{X}_{j}),\ldots,g_{L}(\bm{X}_{j})\}$. Then the dual formulation simplifies to
@@
subject to $\>\>\>\text{no constraint on}\>\>\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma},$
@@
$\quad+\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X}_{i})+\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X}_{i})$
@@
$\quad\times P(\exists\bm{X}_{j}\in R_{j}(\bm{x}^{*}),Z_{j}=1,j=1,\ldots,3^{p})$
@@
Suppose Assumptions Assumption 1: Strong ignorability, either Assumption 2: Positivity or Assumption 2: Relaxed positivity, and Assumption 3: Balance function conditions hold and $\rho$ has a continuous first derivative. Let $\bm{\theta}=(\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma}).$ Then,
@@
We use similar proof techniques to the proof of proposition 1 in Källberg and Waernbaum (2023). First, we prove the following lemma.
@@
$\bm{H}_{k_{1},k_{2}}(\bm{\theta})$ $=E[-e(\bm{X})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X})+\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})c_{k_{2}}(\bm{X})$
$\quad-\{1-e(\bm{X})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X})-\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})c_{k_{2}}(\bm{X})],$
$\bm{H}_{l_{1},l_{2}}(\bm{\theta})$ $=E[-e(\bm{X})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X})+\bm{\gamma}^{T}g(\bm{X})\}g_{l_{1}}(\bm{X})g_{l_{2}}(\bm{X})$
$\quad-\{1-e(\bm{X})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X})\}g_{l_{1}}(\bm{X})g_{l_{2}}(\bm{X})],$
$\bm{H}_{k_{1},l_{1}}(\bm{\theta})$ $=E[-e(\bm{X})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})g_{l_{1}}(\bm{X})$
$\quad-\{1-e(\bm{X})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})g_{l_{1}}(\bm{X})].$
We can construct an estimator for $\bm{H}(\bm{\theta})$, $\hat{\bm{H}}(\bm{\theta})$ using the sample equivalents of $\bm{H}_{k_{1},k_{2}}(\bm{\theta}),\bm{H}_{l_{1},l_{2}}(\bm{\theta}),$ and $\bm{H}_{k_{1},l_{1}}(\bm{\theta})$. Let $c_{i}=-e(\bm{X}_{i})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X}_{i})\}+-\{1-e(\bm{X_{i}})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X_{i}})\}$. By the strict concavity of $\rho$, $c_{i}>0$. Let $v^{*}=[\sqrt{c_{1}},\ldots,\sqrt{c_{n}}]$. Let $v$ be any vector of length $n$. Then $v^{T}\hat{\bm{H}}(\bm{\theta})v=v^{T}\odot v^{*^{T}}B(\bm{X})^{T}B(\bm{X})v\odot v^{*}>0$ since $B(\bm{X})$ is full rank by Assumption Assumption 3: Balance function conditions. Then $\hat{H}$ is positive definite and since $E[\hat{\bm{H}}(\bm{\theta})]=\bm{H}(\bm{\theta})$, $\hat{\bm{H}}(\bm{\theta})\to_{p}\bm{H}(\bm{\theta})$, $\bm{H}(\bm{\theta})$ is positive definite and $l(\bm{\theta})$ is strictly convex as desired. ∎
@@
$\tau$ $=E\left[\frac{Z_{i}Y_{i}}{e(\bm{X}_{i})}\right]-E\left[\frac{(1-Z_{i})Y_{i}}{1-e(\bm{X}_{i})}\right]$
@@
Then, given Proposition 5, Assumption Assumption 3: Balance function conditions, and the assumption that $\rho^{\prime\prime}\{\cdot\}$ is continuous over $\mathbb{R}$, the law of large numbers for averages with estimated parameters applies (Boos and Stefanski (2013) Theorem 7.3) and
@@
Assume $\mu_{1}(\bm{x})=\bm{\beta}_{1}^{T}c(\bm{x})+\bm{\lambda}^{T}g(\bm{x})$ and $\mu_{0}(\bm{x})=\bm{\beta}_{0}^{T}c(\bm{x})+\bm{\lambda}^{T}g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}\in\mathbb{R}^{L}$. We will use a similar proof technique by Zhao and Percival (2017) to show that augmenting the our proposed estimator with estimated outcome regressions does not change the estimator. Let $\hat{\mu}_{0}(\bm{x})$, $\hat{\mu}_{1}(\bm{x})$, and $\hat{e}(\bm{x})$ be outcome regression and propensity score estimates. Then the standard doubly robust estimator for the ATE is,
@@
$\quad+\sum_{Z_{i}=0}w^{*}(\bm{X}_{i})[\hat{\bm{\beta}}_{0}^{T}c(\bm{X}_{i})+\hat{\bm{\lambda}}^{T}g(\bm{X})_{i}]-\frac{1}{n}\sum_{i}\hat{\bm{\beta}}_{0}^{T}c(\bm{X}_{i})+\hat{\bm{\lambda}}^{T}g(\bm{X})_{i}$
@@
$\quad-\hat{\bm{\beta}}_{0}^{T}[\sum_{i}c(\bm{X}_{i})-\sum_{i}w^{*}(\bm{X}_{i})(1-Z_{i})c(\bm{X}_{i})]$
$\quad+\hat{\bm{\lambda}}^{T}[\sum_{i}w^{*}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})-\sum_{i}w^{*}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})]$
@@
and $\hat{\tau}_{DR}=\hat{\tau}_{w^{*}}$. Since $\hat{\tau}_{DR}\to_{p}\tau$ (Bang and Robins, 2005) we also have $\hat{\tau}_{w^{*}}\to_{p}\tau$. ∎
@@
Assume $\mu_{1}(\bm{x})=\bm{\beta}_{1}^{T}c(\bm{x})+\bm{\lambda}_{1}^{T}g(\bm{x})$ and $\mu_{0}(\bm{x})=\bm{\beta}_{0}^{T}c(\bm{x})+\bm{\lambda}_{0}^{T}g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}_{z}\in\mathbb{R}^{L}$. Then, given Proposition 5, Assumption Assumption 3: Balance function conditions, and the assumption that $\rho^{\prime\prime}\{\cdot\}$ is continuous over $\mathbb{R}$, the law of large numbers for averages with estimated parameters applies (Boos and Stefanski (2013) Theorem 7.3) and
@@
$\quad-\frac{1}{n}\sum_{i=1}^{n}w^{*}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}-\mu_{0}(\bm{X}_{i})\}+\frac{1}{n}\sum_{i=1}^{n}\{w^{*}_{0}(\bm{X}_{i})(1-Z_{i})-1\}\mu_{0}(\bm{X}_{i})$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\tau(\bm{X}_{i})-\tau$
$=\frac{1}{n}\sum_{i=1}^{n}S_{i}+R_{1}+R_{2},$
where
@@
$\quad+\sum_{i=1}^{n}\{w_{1}^{*}(\bm{X}_{i})Z_{i}-w_{0}^{*}(\bm{X}_{i})(1-Z_{i})\}\bm{\lambda}^{T}g(\bm{X}_{i}).$
Note that $R_{2}=0$ by the constraints in the optimization problem. Then, since $S_{i}$ takes the same form as the efficient score for the ATE in Hahn (1998), $\hat{\tau}_{w^{*}}$ is asymptotically normal and semiparametrically efficient as long as $R_{1}$ is $o_{p}(n^{-1/2}).$
We will focus on the first component of $R_{1}$, which we decompose as $R_{1}=R_{1}^{1}+R_{1}^{2}$. We can bound the first term as
$|R_{1}^{1}|$ $=\left\lvert\frac{1}{n}\sum_{i=1}^{n}\left(w^{*}_{1}(\bm{X}_{i})-\frac{1}{e(\bm{X}_{i})}\right)Z_{i}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}\right\rvert$
$=\left\lvert\int\left(w^{*}_{1}(\bm{X}_{i})-\frac{1}{e(\bm{X}_{i})}\right)\{Z_{i}Y_{i}-Z_{i}\mu_{1}(\bm{X}_{i})\}dF_{n}(\bm{x})\right\rvert$
$\leq\left\lVert w^{*}_{1}(\bm{X}_{i})-\frac{1}{e(\bm{X}_{i})}\right\rVert\cdot\left\lVert\{Z_{i}Y_{i}-Z_{i}\mu_{1}(\bm{X}_{i})\}\right\rVert$
$=o_{p}(1)O_{p}(n^{-1/2})=o_{p}(n^{-1/2}),$
@@
$\text{Var}(S_{i}+R_{1i})=\text{Var}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}-\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}+\{\tau(\bm{X}_{i})-\tau\}].$
@@
$\quad+\text{Var}\{\tau(\bm{X}_{i})-\tau\}$
$\quad-2\text{Cov}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\},\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}]$
$\quad+2\text{Cov}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\},\{\tau(\bm{X}_{i})-\tau\}]$
$\quad-2\text{Cov}[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\},\{\tau(\bm{X}_{i})-\tau\}].$
We have that
@@
$=-2E[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}]$
$\quad\times E[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}]]$
@@
$\quad\times E[\tilde{w}_{0}(\bm{X}_{i})\{1-e(\bm{X}_{i})\}\{\mu_{0}(\bm{X}_{i})-\mu_{0}(\bm{X}_{i})\}]$
$=0\times 0=0.$